A city is spending $ 19.2 million on a new sewage system. The expected life of the system is 45 years, and it will have no market value at the end of its life. Operating and maintenance expenses for the system are projected to average $ 0.6 million per year. If the city's MARR is 7% per year, What is the capitalized worth of the system? The study period is 100. (Using CW method)
Answer #1

AW = -19200000*(A/P,7%,45)-600000

= -19200000*(0.0735)-600000

= -1411200-600000

= -2011200

CW = -2011200/0.07

= -28731428.57